Solve (𝐷^2 βˆ’ 3𝐷 + 2)𝑦 = π‘₯^2 + sin π‘₯ where 𝐷 =𝑑/𝑑π‘₯

1
Expert's answer
2022-01-02T17:30:49-0500

Corresponding homogeneous diffrential equaion


"(\ud835\udc37^2 \u2212 3\ud835\udc37 + 2)\ud835\udc66 = 0"

Auxiliary equation


"r^2-3r+2=0"

"r_1=1, r_2=2"

The general solution of the homogeneous differential equaion


"y_h=c_1e^x+c_2e^{2x}"

The particular solution of the nonhomogeneous differential equaion


"y_1=Ax^2+Bx+C+D\sin x+E\cos x"

"y_1'=2Ax+B+D\cos x-E\sin x"

"y_1''=2A-D\sin x-E\cos x"

Substitute


"2A-D\sin x-E\cos x"

"-6Ax-3B-3D\cos x+3E\sin x"

"+2Ax^2+2Bx+2C+2D\sin x+2E\cos x"

"=x^2+\sin x"

"2A=1=>A=\dfrac{1}{2}"

"-3+2B=0=>B=\dfrac{3}{2}"

"1-\dfrac{9}{2}+2C=0=>C=\dfrac{7}{4}"

"-3D+E=0"

"D+3E=1"

"D=\dfrac{1}{10}, E=\dfrac{3}{10}"

The general solution of the nonhomogeneous differential equaion


"y=c_1e^x+c_2e^{2x}+\dfrac{1}{2}x^2+\dfrac{3}{2}x+\dfrac{7}{4}""+\dfrac{1}{10}\sin x+\dfrac{3}{10}\cos x"